2. The plate of a parallel plate capacitor, 5.0 x 10 -3 m apart are maintained at a potential
difference of 5.0 x 10 4 V. calculate the magnitude of the:
(i) Electric field intensity between the plate
(ii) Force on the electron
(iii) Acceleration of the electron
(Electronic charge = -1.6 x 10 -19 C, mass of electron = 9.1 x 10 -31 kg)
1
Expert's answer
2020-10-12T07:49:23-0400

1) Since the potential "\phi =\frac W q" , where "W" is energy and "q" is charge, the voltage "U=∆\phi=\frac A q" , where "A" is work, a work is equal to "A=q\times E \times ∆r" , from this follows that

"E=\frac {∆\phi}{∆r}" "= \frac {U}{∆r}=" "\frac {5\times 10^4}{5\times 10^{-3}}=10^7" V.

2) Since "E = \frac {F} {q}" , where "F" is the force, "F = E\times q = 10 ^ 7\times 1 .6\times 10 ^ {-19}" "= 1 .6\times 10 ^ {-12}" N.

3) According to Newton's first law, "F = m\times a" , therefore "a =\frac F m =" "\frac {1 .6\times 10 ^ {-12}} {9 .1\times 10 ^ {-31} }" "\approx 1 .76\times 10 ^ {18}" m/s "^ 2".